OpenStudy (anonymous):

A buffer is created by combining 160 mL of 0.20 M HCHO2 with 75 mL of 0.25 M NaOH. Determine the pH of the buffer. Someone help me to figure out this question..

OpenStudy (rogue):

\[n = MV\]\[n_{HCHO_2} = .160 L * 0.20 M = 0.032 mol \]\[n_{NaOH} = 0.075 mL * 0.25 M = 0.01875mol\]\[OH^- + HCHO_2 \rightarrow H_2O + CHO _{2}^{-}\] 0.01875 moles of formic acid will be converted to formate ions. 0.01875 mol HCHO2 ---> 0.01875 mol CHO2- 0.032 mol - 0.01875 mol = 0.01325 mol HCHO2 \[pH = pK_a - \log_{10} \frac {[HA]}{[A^{-}]}\]You can use moles of HA/A- instead of their concentrations since their molarity and mole ratio are the same. The pKa of formic acid is 3.77.\[pH = 3.77 - \log_{10} \frac {0.01325 mol}{0.01875 mol}\]\[pH = 3.92\]

OpenStudy (rogue):

Do you know if the answer is correct?